find the term 0F AN AP
18,151/2,13..............-47

Grade:11

1 Answers

Arun
6 years ago
a = 18
d = 31/2  – 18= 15.5 –18
 =  –2.5
now
 –47 = 18 + (n –1)*  – 2.5
 –47 –18 =  –2.5 * (n –1)
 –65 / –2.5 = n –1
26 = n –1
n = 27
 
Regards
Arun (askIITians forum expert)